# Short Paper: Atomic Execution is Not Enough for Arbitrage Profit Extraction in Shared Sequencers

### v1

@@ 2.1 Preliminaries
We further assume that the arbitrageur maintains liquidity on both rollups, which is kept in the target tokens $X$ and $Y$. Concretely, the arbitrageur’s liquidity is $L^{X}=L_{A}^{X}+L_{B}^{X}$ (for token $X$) and $L^{Y}=L_{A}^{Y}+L_{B}^{Y}$ (for token $Y$). We can value the total liquidity of the arbitrageur in units of token $Y$ using an external price $P_{\text{ext}}$, and thus, $L=L^{Y}+L^{X}\cdot P_{\text{ext}}$.
@@ 2.3 Trade Sizes
As for $\Delta y_{B}$, the optimal trade occurs when the prices in both pools at the end of the trade are equal, i.e., $P_{A}^{\text{end}}=P_{B}^{\text{end}}$. We can use this to derive $\Delta y_{B}$:
$P_{A}^{\text{end}}=P_{B}^{\text{end}}\Longleftrightarrow$
$\Longleftrightarrow\frac{x_{A}\cdot y_{A}}{\left[x_{A}+(1-f)\Delta x_{A}\right]^{2}}=\frac{\left[y_{B}+(1-f)\Delta y_{B}\right]^{2}}{x_{B}\cdot y_{B}}$
$\Longleftrightarrow\frac{\sqrt{x_{A}\cdot y_{A}}}{x_{A}+(1-f)\Delta x_{A}}=\frac{y_{B}+(1-f)\Delta y_{B}}{\sqrt{x_{B}\cdot y_{B}}}$
$\Longleftrightarrow\frac{\sqrt{x_{A}\cdot y_{A}}}{x_{A}+(1-f)\frac{x_{B}(1-f)\Delta y_{B}}{y_{B}+(1-f)\Delta y_{B}}}=\frac{y_{B}+(1-f)\Delta y_{B}}{\sqrt{x_{B}\cdot y_{B}}}$
$\Longleftrightarrow\frac{\sqrt{x_{A}\cdot y_{A}}}{x_{A}\left[y_{B}+(1-f)\Delta y_{B}\right]+(1-f)x_{B}(1-f)\Delta y_{B}}=\frac{1}{\sqrt{x_{B}\cdot y_{B}}}$
$\Longleftrightarrow\frac{\sqrt{x_{A}\cdot y_{A}}}{x_{A}\cdot y_{B}+x_{A}(1-f)\Delta y_{B}+x_{B}(1-f)^{2}\Delta y_{B}}=\frac{1}{\sqrt{x_{B}\cdot y_{B}}}$
$\Longleftrightarrow x_{A}\cdot y_{B}+x_{A}(1-f)\Delta y_{B}+x_{B}(1-f)^{2}\Delta y_{B}=\sqrt{x_{A}\cdot y_{A}\cdot x_{B}\cdot y_{B}}$
$\Longleftrightarrow\Delta y_{B}=\frac{\sqrt{x_{A}\cdot y_{A}\cdot x_{B}\cdot y_{B}}-x_{A}\cdot y_{B}}{(1-f)x_{A}+(1-f)^{2}x_{B}}$ (10)
@@ 3 Atomicity Profit Conditions
Intuitively, this relationship makes sense. When the external price has a larger difference to the price on rollup $B$ than the price differences in the two rollups, it would be better to arbitrage rollup $B$ against the external price than to arbitrage it against rollup $A$. Therefore, having swap $S_{A}$ failing and swap $S_{B}$ executing leads to a net gain for the arbitrageur when valued against this external price (which means that atomicity is worse for the arbitrageur).
On the other hand, when the external price is smaller than both pool prices (Figure 1(c)), the relationship is inverted. In this case, the rationale is similar, and switching to an atomic regime is only advantageous when failures are more likely on rollup $A$ since arbitraging rollup $A$ against the external price generates more profit than arbitraging it against rollup $B$.
Finally, there is the case where the external price is between the two pool prices (Figure 1(b)). This can arguably be the most likely case if we assume that the external price comes from a more liquid venue and that the pools on the two rollups hover around this more stable price. Interestingly, the expected profit difference is always negative in this case, independently of the failure probabilities. Similarly to the previous cases, when we value liquidity using an external price, and one of the swaps fails and the other executes, we are, in a way, arbitraging the pool that did not fail against the external price. When the external is between the prices in each pool, having only one swap failing is always better than having both reverting, as we would collect some additional profit from arbitraging the pool that did not fail against the external price.
Focusing on the failing probabilities, there is a special case we can analyze. If the two failure probabilities are equal (i.e., $f_{A}=f_{B}$), the expected profit difference is always negative, meaning that, on average, the arbitrageur should stick with the status quo without atomicity. This result comes directly from Equation 12:
$\mathbb{E}[\text{Profit}_{\text{diff}}]=$
$=\Delta x_{B}\big{[}f_{A}(P^{*}_{B}-P_{\text{ext}})+f_{A}(P_{\text{ext}}-P^{*}_{A})+f_{A}^{2}(P^{*}_{A}-P^{*}_{B})\big{]}=$
$=\Delta x_{B}\big{[}f_{A}(P^{*}_{B}-P^{*}_{A})+f_{A}^{2}(P^{*}_{A}-P^{*}_{B})\big{]}=$
$=\Delta x_{B}\cdot f_{A}\cdot(1-f_{A})\cdot(P^{*}_{B}-P^{*}_{A})<0$ (13)

### v2

@@ 2.1 Preliminaries
We further assume that the arbitrageur maintains liquidity on both rollups, which is kept in the target tokens $X$ and $Y$. Concretely, the arbitrageur’s liquidity is $L^{X}=L_{A}^{X}+L_{B}^{X}$ (for token $X$) and $L^{Y}=L_{A}^{Y}+L_{B}^{Y}$ (for token $Y$). We can value the total liquidity of the arbitrageur in units of token $Y$ using an external price $P_{\text{ext}}$, and thus, $L=L^{Y}+L^{X}\cdot P_{\text{ext}}$. Note that $P_{\text{ext}}$ is a theoretical price representing how the arbitrageur values its liquidity. It can be thought of as the price coming from an external source (e.g. an exchange or an oracle) or the price the arbitrage experiences when they settle their liquidity in a future time.
@@ 2.3 Trade Sizes
As for $\Delta y_{B}$, the optimal trade occurs when the prices (excluding fees) in both pools at the end of the trade are equal, i.e., $(1-f)^{2}P_{A}^{\text{end}}=P_{B}^{\text{end}}$. We can use this to derive $\Delta y_{B}$:
$(1-f)^{2}P_{A}^{\text{end}}=P_{B}^{\text{end}}\Longleftrightarrow$
$\Longleftrightarrow(1-f)^{2}\frac{x_{A}y_{A}}{\left[x_{A}+(1-f)\Delta x_{A}\right]^{2}}=\frac{\left[y_{B}+(1-f)\Delta y_{B}\right]^{2}}{x_{B}y_{B}}$
$\Longleftrightarrow\frac{(1-f)\sqrt{x_{A}y_{A}}}{x_{A}+(1-f)\Delta x_{A}}=\frac{y_{B}+(1-f)\Delta y_{B}}{\sqrt{x_{B}y_{B}}}$
$\Longleftrightarrow\frac{(1-f)\sqrt{x_{A}y_{A}}}{x_{A}+(1-f)\frac{x_{B}(1-f)\Delta y_{B}}{y_{B}+(1-f)\Delta y_{B}}}=\frac{y_{B}+(1-f)\Delta y_{B}}{\sqrt{x_{B}\cdot y_{B}}}$
$\Longleftrightarrow\frac{(1-f)\sqrt{x_{A}\cdot y_{A}}}{x_{A}\left[y_{B}+(1-f)\Delta y_{B}\right]+(1-f)x_{B}(1-f)\Delta y_{B}}=\frac{1}{\sqrt{x_{B}\cdot y_{B}}}$
$\Longleftrightarrow\frac{(1-f)\sqrt{x_{A}\cdot y_{A}}}{x_{A}\cdot y_{B}+x_{A}(1-f)\Delta y_{B}+x_{B}(1-f)^{2}\Delta y_{B}}=\frac{1}{\sqrt{x_{B}\cdot y_{B}}}$
$\Longleftrightarrow x_{A}\cdot y_{B}+x_{A}(1-f)\Delta y_{B}+x_{B}(1-f)^{2}\Delta y_{B}=(1-f)\sqrt{x_{A}\cdot y_{A}\cdot x_{B}\cdot y_{B}}$
$\Longleftrightarrow\Delta y_{B}=\frac{(1-f)\sqrt{x_{A}\cdot y_{A}\cdot x_{B}\cdot y_{B}}-x_{A}\cdot y_{B}}{(1-f)x_{A}+(1-f)^{2}x_{B}}$ (10)
@@ 3 Atomicity Profit Conditions
We should highlight that this formula does not depend on the exact design of the Automated Market Maker (AMM). To ease the implementation of the entire simulation , we assumed it was a CPMMas it made the derivation of the optimal trade sizes simpler. However, this formula would also hold for more complex designs such as Uniswap v3.
@@
Intuitively, this relationship makes sense. When the external price has a larger difference to the price on rollup $B$ than the price differences in the two rollups, it would be better to only execute the swap on rollup $B$ than to arbitrage it against rollup $A$. Therefore, having swap $S_{A}$ failing and swap $S_{B}$ executing leads to a net gain for the arbitrageur when valued against this external price (which means that atomicity is worse for the arbitrageur).
On the other hand, when the external price is smaller than both pool prices (Figure 1(c)), the relationship is inverted. In this case, the rationale is similar, and switching to an atomic regime is only advantageous when failures are more likely on rollup $A$ since simply swaping on rollup $A$ generates more profit than arbitraging it against rollup $B$.
Finally, there is the case where the external price is between the two pool prices (Figure 1(b)). Interestingly, the expected profit difference is always negative in this case, independently of the failure probabilities. Similarly to the previous cases, when we value liquidity using an external price, and one of the swaps fails and the other executes, we are, in a way, arbitraging the pool that did not fail against the external price. When the external is between the prices in each pool, having only one swap failing is always better than having both reverting, as we would collect some additional profit from arbitraging the pool that did not fail against the external price.